Linux Meta-FAQ

Michael K. Johnson, johnsonm@redhat.comm v4.6, 20 December 1996

This is the Meta-FAQ for Linux. It is mainly a list of valuable sources of information. Check these sources out if you want to learn more about Linux, or have problems and need help. Lars Wirzenius (wirzeniu@cc.helsinki.fi) wrote the first version of this document, and it is now maintained by Michael K. Johnson (johnsonm@redhat.com). Mail me if you have any questions about this document.

1 Introduction

What is Linux?

Linux is an independent implementation of the POSIX operating system specification, with SYSV and BSD extensions, that has been written entirely from scratch (this means it looks and acts just like Unix). It has no proprietary code in it. Linux is freely distributable under the GNU Public License. It works on IBM PC compatibles with an ISA or EISA bus (including local bus variants VLB and PCI) and a 386 or higher processor. In addition, it also runs on some Amiga and Atari computers with MMU’s. This means 68020 with an external MMU, 68030, 68040, or 68060. See the /pub/linux/680x0 directory at tsx-11.mit.edu. Support for the Digital Alpha is now stable. Red Hat and Craftworks have Alpha distributions of Linux. Support for Sparc is stable, and Red Hat Linux is available for Sparc. Support for PowerPC is in development for multiple platforms, including Nubus and PCI Macintosh, Motorola Powerstack, IBM 830 and 850, and other platforms. Support for ARM and MIPS is in various stages of completion, but don’t hold your breath. Read comp.os.linux.announce instead. See the FAQ for more exact hardware requirements. The Linux kernel is written by Linus Torvalds (Linus.Torvalds@.Helsinki.FI) from Finland, and by other volunteers. Most of the programs running under Linux are generic Unix freeware, many of them from the GNU project.

The Linux FAQ A collection of common problems and their solutions. Answers many questions faster than the net. Stored on many Linux ftp sites (including /pub/linux/docs on tsx-11.mit.edu) and rtfm.mit.edu, the general archive site for all FAQs.

The Linux HOWTO’s These are somewhat like FAQ’s, but instead of answering common questions, they explain how to do common tasks, like ordering a release of Linux, setting up print services under Linux, setting up a basic UUCP feed, etc. See sunsite.unc.edu, directory /pub/Linux/docs/HOWTO/ for the definitive versions of all the HOWTO’s. Other sites with up-to-date copies of the HOWTOs are ftp.cc.gatech.edu and tsx-11.mit.edu.

In addition, there are many short, free-form documents called ”mini-HOWTOs”. These documents cover very specific subjects, such as BogoMIPS or Color-ls. These are available in the directory /pub/Linux/docs/HOWTO/mini/

Linux newsgroups There are several Usenet newsgroups for Linux. It is a good idea to follow at least comp.os.linux.announce if you use Linux. Comp.os.linux.announce is moderated by Lars Wirzenius. To make submissions to the newsgroup, send mail to linux-announce@news.ornl.gov. You may direct questions about comp.os.linux.announce to Lars Wirzenius, wirzeniu@kruuna.helsinki.fi The newsgroup comp.os.linux.announce is a moderated newsgroup for announcements about Linux (new programs, bug fixes, etc). The newsgroup comp.os.linux.answers is a moderated newsgroup to which the Linux FAQ, HOWTO documents, and other documentation postings are made. The newsgroup comp.os.linux.setup is an unmoderated newsgroup for discussion of issues and problems involved in setting up Linux systems. The newsgroup comp.os.linux.admin is an unmoderated newsgroup for discussion of administration of Linux systems. The newsgroup comp.os.linux.development.system is an unmoderated newsgroup specifically for discussion of Linux kernel development. The only application development questions that should be discussed here are those that are intimately associated with the kernel. All other development questions are probably generic Unix development questions and should be directed to a comp.unix group instead, unless they are very Linux-specific applications questions, in which case they should be directed at comp.os.linux.development.apps. The newsgroup comp.os.linux.development.apps is an unmoderated newsgroup specifically for discussion of Linux-related applications development. It is not for discussion of where to get applications for Linux, nor a discussion forum for those who would like to see applications for Linux. The newsgroup comp.os.linux.hardware is for Linux-specific hardware questions. The newsgroup comp.os.linux.networking is for Linux-specific networking development and setup questions. The newsgroup comp.os.linux.x is for Linux-specific X Windows questions. The newsgroup comp.os.linux.misc is an unmoderated newsgroup for any Linux discussion that doesn’t belong anywhere else. In general, do not crosspost between the Linux newsgroups. The only crossposting that is appropriate is an occasional posting between one unmoderated group and comp.os.linux.announce. The whole point of splitting comp.os.linux into many groups is to reduce traffic in each. Those that do not follow this rule will be flamed without mercy...

Other newsgroups Do not assume that all your questions are appropriate for a Linux newsgroup just because you are running Linux. Is your question really about shell programming under any unix or unix clone? Then ask in comp.unix.shell. Is it about GNU Emacs? Then try asking in gnu.emacs.help. Also, if you don’t know another group to ask in, but think there might be, politely ask in your post if there is another group that would be more appropriate for your question. At least the groups comp.unix.{questions,shell,programming,bsd,admin}, and comp.windows.x.i386unix should be useful for a Linux user.

The World-Wide Web Greg Hankins, gregh@cc.gatech.edu, maintains the home WWW page for the Linux project. The URL is http://sunsite.unc.edu/LDP/linux.html

Linux Journal A magazine called Linux Journal was launched several years ago. It includes articles intended for almost all skill levels, and is intended to be helpful to all Linux users. Subscriptions are $22 in the U.S., $27 in Canada and Mexico, and $32 elsewhere around the world, all payable in U.S. funds. Subscription inquiries can be sent via email to subs@ssc.com or faxed to (U.S.) 1-206-782-7191 or mailed to Linux Journal, PO Box 85867, Seattle, WA 98145-1867 USA. SSC has a PGP public key if you wish to send your credit card number via encrypted email: finger info@ssc.com

The Maps Information on free software available for Linux can be found in the Linux Software Map, which can be found on sunsite.unc.edu in /pub/Linux/docs/LSM.gz Information on non-commercial projects can be found in the Projects Map, which is posted occasionally to comp.os.linux.announce and is stored at sunsite.unc.edu in /pub/Linux/docs/Projects-Map/Projects-Map.gz. Information on commercial products can be found in the Linux Commercial Products Map, which is posted occasionally to comp.os.linux.announce

2 Getting Linux

3 Linux FTP sites

A more complete list of Linux FTP sites is in the Linux INFO-SHEET, which can always be found at tsx-11.mit.edu in the directory /pub/linux/docs/. A DVI version is in INFO-SHEET.dvi, and a PostScript version is in INFO-SHEET.ps. The most important sites are listed here; please see the INFO-SHEET for a site nearer to you (there are many mirrors).

textual name             numeric addr    Linux directory
=======================  ==============  ===============
tsx-11.mit.edu           18.86.0.44      /pub/linux
sunsite.unc.edu          152.2.22.81     /pub/Linux
ftp.funet.fi             128.214.248.6   /pub/Linux

These sites are the main “home” sites for Linux where most uploads take place. There are many mirror sites; please use the closest (network-wise) site to you.

3.1 Linux on BBS’s

Occasionally, someone posts a list of BBS’s that have Linux available for download. Try them if you can’t FTP. This list is available by ftp from tsx-11.mit.edu as /pub/linux/docs/bbs.list. You may use ftpmail (described elsewhere in this document) to get this list if you have mail access, or ask a friend with internet access to get it for you.

3.2 Linux on physical media

Linux is distributed on physical media, mainly CD-ROM, by several commercial vendors. Please read the distribution HOWTO, posted regularily to comp.os.linux.announce, and available for ftp at sunsite.unc.edu as /pub/Linux/docs/HOWTO/distribution-HOWTO.

3.3 AFS

Linux is available over AFS by mounting the volume project.linux from sipb.mit.edu

3.4 Commercial networks

GEnie mirrors most of tsx-11.mit.edu and sunsite.unc.edu. Compu$erve also has some Linux archives.

3.5 Mailservers and such

Sunsite offers ftp-mail service — mail ftpmail@sunsite.unc.edu.

4 Linux distributions

Linux is distributed by its author only as a kernel. Other people have put together “distributions” that can be used. These distributions pair the kernel as released by the author with software, to make a complete working package. Most releases include application programs as well as system software, providing “one stop shopping” for Linux.

There are several distributions of Linux, which are available at various sites. Sunsite mirrors almost all the distributions in /pub/Linux/distributions. The most commonly-recommended freely-available distributions are Red Hat, Debian, and Slackware. These are all available for free over the internet, and are also sold on floppies and CD-ROM’s.

There are other distributions of Linux as well. Most commercial distributors of Linux advertise in Linux Journal as well.

5 Linux mailing-lists

Used mostly for discussion between developers of new features and testers of pre-release versions. See addresses in the FAQ. Send mail to majordomo@vger.rutgers.edu with the single word help in the body of the message , and you will get mail explaining how to subscribe to the many Linux mailing lists there. Save this mail, as it tells you how to unsubscribe from the lists, and if you post annoying messages to the list complaining about not being able to get off the list (because you didn’t follow instructions and save the mail telling you how to unsubscribe), you will likely be flamed for wasting international bandwidth and money.

6 Documentation for various programs

Many programs come with some sort of documentation, often in a file called README or something similar. It is a VERY good idea to read them with care. It is boring to see (and answer) questions that are answered in the documentation.

7 More Documentation

The Linux Documentation Project is working on a lot of documentation. Already, over 1600 pages of book-style documentation has been released to the general public, and a large set of man pages has also been released, with more to follow. Check sunsite.unc.edu:/pub/Linux/docs/LDP for documents written by the LDP. The LDP also provides the HOWTO’s, described above.

8 Keeping track of current releases

Important new releases, programs, and ports are usually announced in comp.os.linux.announce. finger

torvalds@klaava.helsinki.fi or finger @linux.cs.helsinki.fi to get some information about the

current kernel.

9 This Document

The latest version of this document should always be available from

<http://sunsite.unc.edu/LDP/linux.html>

10 Legalese

Trademarks are owned by their owners. Satisfaction not guaranteed. No warranties about this document. Void where prohibited.

10. Legalese 6 The content of this document is placed in the public domain, but if you quote it, please be polite and attribute your source.